# Can Large Language Models Detect Misinformation in Scientific News Reporting?

### v1

@@ 4.2.7. Prompt Strategies
  • Zero-shot prompting: LLMs are presented with a task without any prior specific training or examples related to that task (Brown et al., 2020). We test the performance of LLM on scientific misinformation detection by zero-shot prompts directly using the pre-existing knowledge of the LLMs.
@@
  • Few-shot prompting: In contrast to zero-shot prompting, where LLMs are presented with tasks without prior training or examples, few-shot prompting involves furnishing LLMs with a concise set of examples prior to task execution (Brown et al., 2020). This approach is designed to provide the model with essential context, thereby augmenting its capability for tasks like detecting scientific misinformation. In our methodology, we supply the LLMs with two carefully selected examples: one deemed ’reliable’ with accompanying reasoning, and another labeled ’unreliable’. This strategy is intended to better equip the LLMs to discern and categorize information accurately in the execution of the task.
@@
  • Chain-of-thought prompting (CoT): CoT prompting involves structuring prompts to elicit a step-by-step reasoning process, effectively emulating the cognitive process humans employ in solving complex problems (Xu et al., 2022). In our approach, we used the dimensions of scientific validity defined in Section 4.1 to design multiple CoT prompts to guide the LLMs. This methodology not only aids the LLMs in systematically dissecting and assessing factual content but also aligns their reasoning process with structured, human-like analytical methods.
5. Results and Discussions
@@
We used GPT-3.5, GPT-4, Llama2-7B, and Llama2-13B in our architectures with the temperature set to 00. This setting ensures that the LLMs generate responses with the highest predictability, thereby maintaining uniformity across the experiments. The performance of each of the proposed architectures, using each of the above LLMs is measured using accuracy, precision, recall, and F1-score. These results are displayed in Tables  2 and 3.
We note immediately that there is a significant difference between the performances of GPT and Llama2, with both GPT-3.5 and GPT-4 beating Llama2-7B/13B. All Llama2-7B/13B models achieved an accuracy score barely above random guessing. This may be due to the greater popularity of GPT, which may have made the crowdsourced, ongoing training of the GPT models stronger than those of Llama2.
@@ 5.1. Human-Written vs LLM-Generated Misinformation
Table 2 records the results of our experiments on the SERIf architecture and Table 3 shows the performance of the SIf and D2I architectures. From these tables, we note that there is a significant difference in misinformation detection accuracy and F1 scores for the human-written articles and LLM-generated articles. Across all architectures, it is consistently more challenging to identify LLM-generated scientific misinformation compared to human-authored misinformation. This is evidenced by high recall scores paired with low precision scores, indicating a tendency of the detectors to classify news as True. Such a trend highlights the difficulty in discerning false information within LLM-generated scientific news. This notable discrepancy emphasizes the complexity and challenge inherent in detecting misinformation in LLM-generated content. It aligns with similar findings in non-scientific misinformation domains (Chen and Shu, 2023). These results also raise concerns about the potential misuse of LLMs and underscore the importance of advancing our detection methodologies to keep pace with the evolving capabilities of LLMs, considering their implications for public safety.
Table 3. Performance results of our proposed SIf and D2I architectures using different LLMs and prompt strategies.
Models Arch Prompt Strategy Human-Written LLM-Generated Overall
@@
GPT3.5 SIf Zero-Shot 82.50 80.00 86.50 83.50 62.00 57.00 99.00 72.00 72.25 68.50 89.13 77.75
Few-Shot 78.67 79.76 76.83 78.27 65.33 59.89 92.83 72.81 72.00 69.83 84.83 75.54
CoT 76.08 79.88 73.67 75.68 70.17 65.12 86.83 74.43 73.13 72.50 80.25 75.06
CoT 77.17 69.50 96.83 80.91 64.08 57.63 99.65 73.03 70.63 63.57 98.24 76.97
GPT-4 SIf Zero-Shot 80.00 80.00 79.00 80.00 71.00 64.00 98.00 77.00 75.50 72.00 88.50 78.50
CoT 73.92 82.38 60.83 69.98 71.08 65.79 87.83 75.23 72.50 74.09 74.33 72.61
CoT 78.08 84.60 68.67 75.80 72.00 65.20 98.50 78.30 75.04 74.90 83.56 77.05
@@ 5.2. Comparison Across Architectures (RQ1)
The comparative analysis revealed a significant performance superiority of GPT over Llama2. Therefore, GPT was selected for further evaluation of the various architectures proposed in this work. According to the result presented in Table 3, it is evident that the SIf architecture yields the highest accuracy in processing human-authored articles, achieving an impressive $82.50\%$ in accuracy and $83.50\%$ in F1 score. The encouraging results, even in the absence of the ‘sentence-level evidence retrieval’ module, suggest the potential to develop more flexible and generalized scientific misinformation detection models. In stark contrast, the performances of the D2I models were notably subpar. This disparity indicates that merely supplying a Large Language Model (LLM) with a news article coupled with scientific articles does not suffice to guarantee effective performance. Furthermore, our results highlight the importance of the ‘summarization’ module. By distilling key statements from the news, this module minimizes the impact of extraneous information, thereby enhancing the LLM’s ability to generate more accurate predictions.
Table 4. Example of prompt and response for generating explanations.
Prompt: Decide if the evidence ’support’ or ’refute’ the point of view of the news paragraph. Provide your prediction and explain your reasoning in a few sentences.
Response: Refute. The evidence provided does not support the claim made in the news paragraph. While the evidence mentions neurological complaints reported by vaccinated individuals, they do not provide any evidence that suggests COVID-19 mRNA vaccines cause extensive brain damage. Additionally, the fact that different vaccines produced different profiles of complications does not necessarily indicate brain damage. The study’s authors calling for further investigations does not confirm the claim either.
5.3. Comparison Across Prompting Strategies (RQ2)
Observing the overall performance from Table 2 and 3, the results reveal a significant trend: the CoT prompting generally outperforms the zero-shot approach. Notably, for LLM-generated data, the CoT prompt markedly enhances detection capabilities. This suggests that our defined dimensions of scientific validity effectively aid LLMs in making more accurate predictions. However, an interesting observation in the SIf architecture for human-authored content is that, while the CoT prompt led to higher precision, it resulted in lower recall compared to the zero-shot approach. This indicates that our scientific validity dimensions sometimes prompt the LLM to apply more stringent criteria in its judgments.
On the other hand, the few-shot setting did not significantly improve performance, implying that despite providing two well-crafted examples (one positive and one negative), it is challenging for LLMs to extract substantial features from the provided cases. This not only highlights the complexity of scientific misinformation detection but also underscores the intricate nature of the potential scientific misinformation data involved.
@@ 5.4. Explainability Study (RQ3)
To assess the explainability of the proposed architectures, we prompt the LLMs to not only classify the news articles as reliable or unreliable, but also to explain the reasoning behind these classifications. An example of the prompt and the result are presented in Table 4. The response illustrates the LLM’s ability to the content of the news article against corresponding evidence and articulates the differences, leading to explanations that are comprehensible to users. This feature is crucial as it aids individuals in discerning the veracity of scientific content, enhancing their understanding of what constitutes true versus false science. We provide the complete example in the appendix C.
@@ 6. Conclusions
[…]
@@ Acknowledgements.
To Robert, for the bagels and explaining CMYK and color spaces.
@@ A.4. Review Guidance for Human-Written Articles
A.5. An Example of LLM-Generated Data Sample
@@
Appendix B A Summary of Prompt Examples
@@
Table 5. An example of the Zero-Shot prompt.
System Message: As a Fact Checker, your role involves analyzing a news paragraph and several evidence sentences provided by the user. The user will present a news paragraph. Following this, the user will present evidence sentences. Your task is to determine the factual accuracy of the news story based on these evidence sentences. To justify your conclusion, select and reference specific phrases or sentences from both the news story and the evidence provided.
User Message: I will give you one news paragraph and several relevant sentences. Please help me determine if these sentences support or refute the news point of view. Finally, please answer using one word ‘refute’ or ‘support’ and give reasons. Please provide the final output in JSON format containing the following two keys: prediction and reason.
Table 6. An example of the Few-Shot prompt.
System Message: As a Fact Checker, your role involves analyzing a news paragraph and several evidence sentences provided by the user. The user will present a news paragraph. Carefully read this paragraph to understand its central claim. Following this, the user will present evidence sentences. These sentences may either support or refute the news paragraph’s central claim. Your task is to determine the factual accuracy of the news story based on these evidence sentences. Are they supporting or contradicting the news? To justify your conclusion, select and reference specific phrases or sentences from both the news story and the evidence provided.
User Message: Task: Analyze the following news paragraph and several relevant sentences to determine their relationship.
Example 1: {One positive example with label and reason.}
Example 2: {One negative example with label and reason.}
Now analyze the following: {News Paragraph} and {Evidence Corpus}
Instructions: Decide if the relevant sentences ’support’ or ’refute’ the point of view of the news paragraph. Provide your answer in one word - either ‘support’ or ‘refute’. Then, explain your reasoning in a few sentences. Output: format your response as JSON with two keys: prediction and reason.
Table 7. An example of the Chain-of-Thought (CoT) prompt involves integrating our defined dimensions of scientific validity into the thought process, guiding the LLM toward making decisions.
System Message: You are a Fact Checker. The user will present a new paragraph. Following this, the user will present evidence paragraphs. These sentences may either support or refute the news paragraph’s central claim. Your task is to determine the factual accuracy of the news story based on these evidence paragraphs. Make a final prediction and provide a comprehensive explanation step by step based on the following:
@@
User Message: I will give you one news paragraph and relevant evidence corpus. Please help me determine if these paragraphs support or refute the news point of view. Finally, please answer using one word ‘refute’ or ‘support’ and give reasons. Please provide the final output in JSON format containing the following two keys: prediction and reason.
Appendix C More Details of Explainability Study
The example in figure 5 is from SERIf architecture under the CoT prompt and GPT-4. It shows how the ’sentence-level evidence retrieval’ module effectively identifies relevant statements from the evidence corpus, detects contradictions between the original text and evidence, and makes accurate predictions during the ’Inference’ phase based on the predefined dimensions of scientific validity. Such effective explanations enhance understanding of the reasoning process.

### v2

@@ 4.2.7. Prompt Strategies
  • Zero-shot prompting: LLMs are presented with a task without any prior specific training or examples related to that task (BroEtal20). We test the performance of LLM on scientific misinformation detection by zero-shot prompts directly using the pre-existing knowledge of the LLMs.
@@
  • Few-shot prompting: In contrast to zero-shot prompting, where LLMs are presented with tasks without prior training or examples, few-shot prompting involves furnishing LLMs with a concise set of examples prior to task execution (BroEtal20). This approach is designed to provide the model with essential context, thereby augmenting its capability for tasks like detecting scientific misinformation. In our methodology, we supply the LLMs with two carefully selected examples: one deemed ’reliable’ with accompanying reasoning, and another labeled ’unreliable’. This strategy is intended to better equip the LLMs to discern and categorize information accurately in the execution of the task.
@@
  • Chain-of-thought prompting (CoT): CoT prompting involves structuring prompts to elicit a step-by-step reasoning process, effectively emulating the cognitive process humans employ in solving complex problems (WeiEtal22). In our approach, we used the dimensions of scientific validity defined in Section 4.1 to design multiple CoT prompts to guide the LLMs. This methodology not only aids the LLMs in systematically dissecting and assessing factual content but also aligns their reasoning process with structured, human-like analytical methods.
We display all Prompts used in the experiment at Appendix D
5. Experiment and Results
5.1. Experiment Setup
Baseline Setup
The CoSMis (SciNews) dataset aims to address a significant gap left by previous datasets, which involved manual claim generation steps while no original articles were provided (such as SciFact (wadden2020fact) and Check-Covid (wang2023check)). This limitation from previous works makes it challenging to directly apply these datasets within our framework. Despite these challenges, we have established a baseline using BERT-based models to enhance our analytical rigor. We treat it as an Natural Language Reasoning (NLR) task. Given a news or summarized news paragram and relevant selected evidence from the evidence corpus, the reasoning model acts as an evaluator to identify a pair of news/summarized news and related evidence as true or false. The model input will be $[News<SEP>EvidenceSentence]$ or $[SummarizedNews<SEP>EvidenceSentence]$. The ‘Summarized News’ and ‘Evidence Sentences’ come from our best-performing experimental step (SIF with Zero-Shot setting by using GPT-4). We choose two pre-trained models as baseline: BERT (devlin2018bert) and SciBERT (beltagy2019scibert). For SciBERT, it trained using masked language modeling on a large corpus of scientific text. We would like to understand how different the models are that include domain information versus those that do not include domain information.
Implementation Details
@@ 5.1. Experiment Setup
We first employed GPT-3.5, GPT-4 with the temperature set to 0 and Llama2-7B/13B/70B, Llama3-8B with the temperature set to $0.0001$ on the SERIf architecture. This setting ensures that the LLMs generate responses with the highest predictability. The performance of each of the proposed architectures, using each of the above LLMs is measured using accuracy, precision, recall, and F1-score. From the results in Table 2, we see that the GPT models perform significantly better than the LLAMA series. All Llama models achieved an accuracy score barely above random guessing. Hence we used the GPT models for all other architectures (SIF+D2I).
The baseline experiment is implemented by using PyTorch. Since the baseline experiment involves a training step, our dataset must be divided into a training set and a test set. To analyze the baseline method’s dependence on training data, we split the dataset using two schemes: a 5:5 ratio and an 8:2 ratio, respectively.
@@ 5.2. Human-Written vs LLM-Generated Misinformation
Table 2 records the results of our experiments on all architectures. From this tables, we note it is consistently more challenging to identify LLM-generated scientific misinformation compared to human-written misinformation, across all architectures. This is evidenced by high recall scores paired with low precision scores, indicating poor True Negative (TN) prediction and a propensity for the detectors to misclassify news as ‘Reliable’. and a tendency of the detectors to classify news as Reliable. Such a trend highlights the difficulty in discerning false information within LLM-generated scientific news. This aligns with similar findings in non-scientific misinformation domains (CheShu23). These results also raise concerns about the potential misuse of LLMs and underscores the importance of advancing our detection methodologies to keep pace with the evolving capabilities of LLMs, considering their implications for public safety.
We further analyzed the detection performance of different LLMs on the CoSMis (SciNews) dataset. From Table 2, the Llama models underperformed significantly, leading us to discontinue their use in further tests of different model structures. Notably, Llama2-70B performance slightly superior to Llama2-13B and Llama3-8B, and 13B/8B also outperformed the 7B model. However, the overall results of Llama2-70B were still underwhelming. It is important to note that the results in Table 2 exhibit a pattern of ‘high recall, low precision’ in both ‘Human-Written’ and ‘LLM-Generated’ categories, indicating that all Llama models readily classifies text as ‘Reliable’. This suggests that Llama may have a limited capacity for distinguishing between nuanced cases, thereby reducing its ability to handle complex reasoning effectively. In contrast, GPT-3.5 (340B) demonstrated significant improvements, with GPT-4 delivering the best performance, indicating a strong correlation between increased model parameters and enhanced reasoning capabilities. Furthermore, this also suggests that GPT models have better reliability when used in real-world scientific misinformation detection scenarios.
Table 3. Performance results of baseline models. ‘N+ES’ denotes ‘News + Evidence Sentence’, ‘SN+ES’ denotes ‘Summarized News + Evidence Sentence’.
Train-Test Ratio Model Input Text Human-Written LLM-Generated Overall
@@
50-50 BERT N+ES 46.17 47.11 62.50 53.67 54.58 52.67 90.33 66.51 50.38 49.89 76.42 60.09
SN+ES 48.42 48.82 65.50 55.90 56.75 53.90 93.67 68.51 52.64 51.36 79.59 62.21
SciBERT N+ES 47.75 48.31 64.17 55.09 58.58 54.85 97.00 72.11 53.17 51.58 80.59 63.60
SN+ES 49.92 49.94 68.17 57.66 62.00 57.00 99.00 72.00 55.96 53.47 83.59 64.83
80-20 BERT N+ES 53.33 53.33 100 69.56 49.44 49.44 100 66.17 51.37 51.37 100 67.87
SN+ES 54.72 54.72 100 70.74 53.33 53.33 100 69.56 54.02 54.02 100 70.17
SciBERT N+ES 68.58 66.92 73.50 70.08 69.44 95.35 44.79 60.99 69.01 81.14 59.15 65.54
SN+ES 71.25 69.59 75.50 72.38 69.75 68.60 72.83 70.63 70.5 69.10 74.17 71.51
@@ 5.3. Comparison Across Architectures (RQ1)
From Table 2 it is evident that the SIf architecture performs best overall with $75.50\%$ accuracy and $78.50\%$ F1 score. The encouraging results, even in the absence of the ‘sentence-level evidence retrieval’ module, suggest the potential to develop more flexible and generalized scientific misinformation detection models. By contrast, the performances of the SERIf and D2I models were notably subpar when zero-shot prompting was used; however, the performance improves significantly, when paired with DoV-guided CoT prompting. This shows that incorporating DoV in CoT prompting can improve performance. Furthermore, our results highlight the importance of the ‘summarization’ module. In the zero-shot setting, the results for DI2 were significantly lower than those for SIf and SERIf. By distilling key statements from the news, this module minimizes the impact of extraneous information, thereby enhancing the LLM’s ability to generate more accurate predictions.
5.4. Comparison Across Strategies (RQ2)
From Table 2, we see a significant trend: the DoV-CoT prompting generally outperforms the zero-shot and few-shot prompting. Notably, for LLM-generated data, the DoV-CoT prompt markedly enhances detection capabilities. This suggests that our proposed dimensions of scientific validity effectively aid LLMs in making more accurate predictions. However, an interesting observation in the few-shot setting is that it did not significantly improve performance, implying that despite providing two well-crafted examples (one positive and one negative), it is challenging for LLMs to extract substantial features from the provided cases. This not only highlights the complexity of scientific misinformation detection but also underscores the intricate nature of the potential scientific misinformation data involved.
@@ 5.5. Explainability Study (RQ3)
To assess the explainability of the proposed architectures, we prompt the LLMs to not only classify the news articles as reliable or unreliable but also to explain the reasoning behind these classifications and score the news article along the dimensions of scientific validity (DoV) using a number in [-1,1]. The examples of the prompt and the result for the SIf architecture using the CoT prompt and GPT-4 are presented in Fig. 4 and Appendix LABEL:appendix_explain, Fig. 9 (last page). In detail, Fig. 4 shows a spider plot of the scores along each DoV. For the “unreliable” example (left), the news paragraph received a score of -1 in Alignment, Causation Confusion, Accuracy, and Generalization, and a score of 0 in Contextual Fidelity. For the ”reliable” example (right), the news paragraph received a score of 1 in Alignment, Accuracy, and Contextual Fidelity, and a score of 0 in Causation Confusion, and Generalization. A spider-plot such as this provides a clear picture of which DoV is violated for any given input scientific news article. In addition, such a spider plot can provides a comprehensive visual representation of the DoV-CoT reasoning results.
Figure 4. Comparison of two spider plot visualizations: The left side corresponds to the ’Unreliable’ case, while the right side corresponds to the ’Reliable’ case. By visualizing the ’axis of scientific validity,’ we can clearly observe the process of the LLM applying DoV to evaluate scientific news and the resulting differences.
5.6. Comparing Baselines:
The Table 3 shows the results of the baseline experiment. Although the experimental results under an 80%:20% data split can be compared with some of LLM pipelines’ results, when the proportion of the training set is reduced to 50%, the overall prediction performance significantly decreases, which is far inferior to that of the LLM pipeline. This indicates that traditional methods are highly dependent on the training set size, which may not be suitable for contemporary real-world scenarios characterized by the daily proliferation of vast amounts of misinformation from different sources, where it reflects the need to establish a detection pipeline using LLMs. Furthermore, Table 3 also indicate that combining evidence with summarized news articles (SN+ES) yields better outcomes than using news and evidence directly (N+ES). This underscores the importance of the “summarization” block and its generalization across different frameworks. Additionally, the domain-specific SciBERT get the reasonable ‘Precision’ and ’Recall’ score under the 80%:20% data split and also outperforms BERT results, highlighting the value of domain knowledge. This analysis motivates further fine-tune the LLMs on the scientific domain corpus to enhance the scientific misinformation detection.
@@ 6. Conclusions
In this paper, we explore LLMs for identifying unreliable scientific news ‘in the wild’. We created the CoSMis (SciNews) dataset, which includes both human-written and LLM-generated articles, each verified against scientific literature. We defined specific dimensions of scientific validity for news misinformation and introduced three LLM-based architectures for identifying unreliable scientific news. Our tests across various LLMs and prompting strategies yielded key insights: 1) DoV-CoT prompting can improve performance in general 2) with appropriately designed pipelines and prompting strategies, LLMs’ offer a viable approach scientific misinformation detection in the wild, since they offer a way to approach this problem without extensive training, 3) in general it is harder to identify LLM-generated misinformation, and 4) LLMs can provide rationales for their judgments.
@@ Acknowledgements.
The authors would like to thank John, Chumki, and David for their assistance in generating some of the LLM-generated data and quality control.
@@ A.4. Review Guidance for Human-Written Articles
Appendix B Statistics of CosMis Dataset
The CosMis is a balanced dataset that contains an equal number of human-written and LLM-generated news articles on each label. We further analyzed the proposed CosMis Dataset:
  • For human-written articles part: maximum number of sentences in an article is 557; minimum number of sentences is 6; The average number of sentences per article is 54.49. The average number of words per sentence within all the news articles is 19.39.
  • For LLM-generated part: maximum number of sentences in an article is 35; minimum number of sentences is 1; The average number of sentences per article is 8.24; and average number of words per sentence within all the news articles: 21.88.
Then, we visualized the distribution of sentence length as well as the average number of sentences in the dataset.
Figure 5. The comparison between the number of sentences in Human-Written Articles and the number of sentences in LLM-Generated Articles .
Figure 6. The number of sentences in LLM-Generated Articles.
In figure 5, it is evident that the human-written article is longer than the LLM-generated article. This discrepancy arises due to the token limit imposed on LLM outputs. Since the input prompt includes the abstract from a scientific paper, a significant portion of the token allocation is consumed, thereby limiting the length of the LLM-generated article. Despite this, the shape of two distributions in figure 5 are remarkably similar. Further analysis of figure 6 reveals a high consistency in the distribution of sentence lengths, suggesting that LLMs are capable of producing articles that closely mimic human writing.
Figure 7. The orange line is the BERT-Score distribution between the scientific abstract and generated true article’ and the blue line is the BERT-Score distribution between abstract and generated false article.
Appendix C Quality Control(QC)
C.1. QC for LLM-generated News Articles
Additionally, we employed the BERT-Score (zhangbertscore) (another commonly used assessment method for text generation), which assesses contextual understanding and lexical polysemy, providing a nuanced evaluation, to evaluate the filtered LLM-generated part again. After calculating, the average BERT-Score between the scientfic abstract and the LLM-generated false article was 0.8269, while ‘Abstract - Generated True Article’ scored higher at 0.9127.
By examining the distributions presented in Figure 7, we observed a significant difference between the two Bert-Scores. This discrepancy not only validates the effectiveness of our filtration process but also confirms the authenticity of the fake articles we generated.
C.2. An Example of LLM-Generated Data Sample
We designed the ‘Jailbreak’ prompt to enable the LLMs to generate fake news-oriented scientific articles that were both informative and contextually aligned with the provided abstracts. The designed jailbreak prompt is illustrated in Figure 2.
Below is the detailed Prompt for the experiment:
  • System Prompt: You are a journalist. The user is developing a course that will help people understand the difference between false and real science when reported in the popular press. The user needs you to complete some article writing based on the content provided by the user.
  • User Prompt: Hi, I am developing a course that will help people understand the difference between false and real science when reported in the popular press. Could you write two newspaper-style articles covering the content based on {the selected scientific abstract}, one True and one that includes false statements but it is very convincing, that students in the class could compare? Label the two versions ‘True’ and ‘Convincing False’.
@@ C.2. An Example of LLM-Generated Data Sample
Appendix D A Summary of Prompt Examples
@@
Zero-Shot Prompt. System Message: As a Fact Checker, your role involves analyzing a news paragraph and several evidence sentences provided by the user. The user will present a news paragraph. Following this, the user will present evidence sentences. Your task is to determine the factual accuracy of the news story based on these evidence sentences. To justify your conclusion, select and reference specific phrases or sentences from both the news story and the evidence provided. User Message: I will give you one news paragraph and several relevant sentences. Please help me determine if these sentences support or refute the news point of view. Finally, please answer using one word ‘refute’ or ‘support’ and give reasons. Please provide the final output in JSON format containing the following two keys: prediction and reason.
Few-Shot Prompt. System Message: As a Fact Checker, your role involves analyzing a news paragraph and several evidence sentences provided by the user. The user will present a news paragraph. Carefully read this paragraph to understand its central claim. Following this, the user will present evidence sentences. These sentences may either support or refute the news paragraph’s central claim. Your task is to determine the factual accuracy of the news story based on these evidence sentences. Are they supporting or contradicting the news? To justify your conclusion, select and reference specific phrases or sentences from both the news story and the evidence provided. User Message: Task: Analyze the following news paragraph and several relevant sentences to determine their relationship. Example 1: {One positive example with label and reason.} Example 2: {One negative example with label and reason.} Now analyze the following: {News Paragraph} and {Evidence Corpus} Instructions: Decide if the relevant sentences ’support’ or ’refute’ the point of view of the news paragraph. Provide your answer in one word - either ‘support’ or ‘refute’. Then, explain your reasoning in a few sentences. Output: format your response as JSON with two keys: prediction and reason.
DoV Chain-of-Thought Prompt. System Message: You are a Fact Checker. The user will present a new paragraph. Following this, the user will present evidence paragraphs. These sentences may either support or refute the news paragraph’s central claim. Your task is to determine the factual accuracy of the news story based on these evidence paragraphs. Make a final prediction and provide a comprehensive explanation step by step based on the following:
@@
User Message: I will give you one news paragraph and relevant evidence corpus. Please help me determine if these paragraphs support or refute the news point of view. Please answer using one word ‘refute’ or ‘support’ and give reasons. Then, score the news article based on each axis of scientific validity between [-1, 1] under the keyword: ’scores’. For scoring, assign a float value in the range -1 and 1 to each axis, where -1 indicates strong disagreement, 0 indicates neutrality, and 1 indicates strong agreement. Please provide the final output in JSON format containing the following three keys: prediction, reason and scores.
Appendix E More Details of Explainability Study
An example prompt and response for the SIf architecture using the CoT prompt and GPT-4 is shown in Fig. 9. The articles used in this example is two human-written articles. It shows how the SIf architecture effectively identifies relevant statements from the evidence corpus, detects contradictions between the original text and evidence, and makes accurate predictions during the ‘Inference’ phase based on the predefined dimensions of scientific validity. Such effective explanations enhance understanding of the reasoning process. Further, we visualized the scores, as shown in Fig. LABEL:fig:spider_comparison. This helps the user to quickly understand whether DoV the news does match or not.